40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11258     Entry time: Mon Apr 27 01:13:08 2015
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRCL angular FF not working, no locking :( 

I'm sad.  And frustrated. 

The PRCL angular feed forward is not working, and without it I am having a very difficult time keeping the PRMI locked while the arms are at high power (either buzzing, or the one time I got stable high power partway through the transition).  Obviously if the PRMI unlocks once CARM and DARM are mostly relying on the REFL signals, I lose the whole IFO. 

Q and I had been noticing over the last few weeks that the angular feed forward wasn't seeming quite as awesome as it did when I first implemented it.  We speculated that this was likely because we had started DC coupling the ITM optical levers, which changes the way seismic motion is propagated to cavity axis motion (since the ITMs are reacting differently).

Anyhow, today it does not work at all.  It just pushes the PRM until the PRMI loses lock. I am worried that, even though Rana re-tuned the BS and PRM oplev servos to be very similar to how they used to be, there is enough of a difference (especially when compounded with the DC coupled ITMs) that the feed forward transfer functions just aren't valid anymore.

Since this prevents whole IFO locking, I spent some time trying to get it back under control, although it's still not working. 

I remeasured the actuator transfer function of how moving PRM affects the sideband spot at the QPD, in the PRMI-only situation.  I didn't make a comparison plot for the yaw degree of freedom, but you can see that the pitch transfer function is pretty different below ~20Hz, which is the whole region that we care about.  In the plot below, black is from January (PRMI-only, no DC-coupled ITMs) and blue is from today (PRMI-only, with DC-coupled ITMs, and somewhat different BS/PRM oplev setup):

Pitch_oldVsNew.pdf

I calculated new Wiener filters, and tried to put them in, but sometimes (and I don't understand what the pattern is yet) I get "error" in the Alternate box, rather than the zpk version of my sos filter.  It seems to go away if you use fewer and fewer poles for fitting the Wiener filters, but then the fit is so poor that you're not going to get any subtraction (according to the residual estimation plot that uses the fitted filters rather than the ideal Wiener filters). The pitch filters could only handle 6 poles, although the yaw filters were fine with 20.

The feed forward just keeps pushing the PRM away though.  I flipped the signs on the Wiener filters, I tried recalculating without the actuator pre-filtering, I don't know why it's failing.  But, I'm not able to lock the interferometer.  Which sucks, because I was hoping to finally get most of my noise coupling measurements done today.

 

Attachment 1: Pitch_oldVsNew.pdf  16 kB  | Hide | Hide all
Pitch_oldVsNew.pdf
ELOG V3.1.3-